a.) Consider the point A=(-1,0,1), B=(0, - 2,3) and C = (-4,4,1) to be vertices of a triangle "\Delta" . Evaluate all side lengths of

"\Delta"

b.) let "\Delta" be the triangle with vertices the points P=(3, 1,-1), Q=(2, 0,3) and R=(1, 1,1). Determine whether "\Delta" is a right triangle. If it is not, explain with reason, why?

c.) let u=<0,1,1>, v=<2,2,0> and w=<-1,1,0> be three vectors in standard form. (i) Determine which two vectors form a right angle triangle? (ii) find "\theta" =uw, the angel between the given two vectors.

d.) let x<0.find the vector n=<x, Y, z> that is orthogonal to all three vectors u=<1,1,-2>,v=<-1,2,0> and w=<-1,0,1>.

e.)find a unit vector that is orthogonal to both u =<0,-1,-1> and v=<1,0,-1>


1
Expert's answer
2021-06-25T05:45:57-0400

a)


"\overrightarrow{AB}=\langle0-(-1),-2-0,3-1 \rangle=\langle1,-2,2 \rangle"


"AB=\sqrt{(1)^2+(-2)^2+(2)^2}=3"


"\overrightarrow{AC}=\langle-4-(-1),4-0,1-1 \rangle=\langle-3,4,0 \rangle"


"AC=\sqrt{(-3)^2+(-4)^2+(0)^2}=5"


"\overrightarrow{BC}=\langle-4-0,4-(-2),1-3 \rangle=\langle-4,6,-2 \rangle"


"BC=\sqrt{(-4)^2+(6)^2+(-2)^2}=2\sqrt{14}"

b)


"\overrightarrow{PQ}=\langle2-3,0-1,3-(-1) \rangle=\langle-1,-1,4 \rangle"


"\overrightarrow{PR}=\langle1-3,1-1,1-(-1) \rangle=\langle-2,0,2 \rangle"

"\overrightarrow{QR}=\langle1-2,1-0,1-3 \rangle=\langle-1,1,-2 \rangle"

"\overrightarrow{PQ}\cdot \overrightarrow{PR}=-1(-2)-1(0)+4(2)=10\not=0"

"\overrightarrow{PQ}\cdot \overrightarrow{QR}=-1(-1)-1(1)+4(-2)=-8\not=0"

"\overrightarrow{QR}\cdot \overrightarrow{PR}=-1(-2)+1(0)-2(2)=-2\not=0"

There is no a pair of orthogonal vectors, hence "\vartriangle" is not right.


c)


"\vec u\cdot\vec v=\langle 0,1,1 \rangle\langle 2,2,0 \rangle=0+2+0=2\not=0"

"\vec u\cdot\vec w=\langle 0,1,1 \rangle\langle -1,1,0 \rangle=0+1+0=1\not=0"

"\vec w\cdot\vec v=\langle -1,1,0 \rangle\langle 2,2,0 \rangle=-2+2+0=0"

"\vec w \perp \vec v"

Vectors "\vec w" and "\vec v" form a right angle triangle.



"\vec u\cdot\vec w=\langle 0,1,1 \rangle\langle -1,1,0 \rangle=0+1+0=1"

"||\vec u||=\sqrt{(0)^2+(1)^2+(1)^2}=\sqrt{2}"

"||\vec w||=\sqrt{(-1)^2+(1)^2+(0)^2}=\sqrt{2}"

"\cos \theta=\dfrac{\vec u\cdot\vec w}{||\vec u||\cdot||\vec w||}=\dfrac{1}{\sqrt{2}\sqrt{2}}=\dfrac{1}{2}"

"\theta=60\degree"

d)


"\vec u\cdot\vec n=\langle 1,1,-2 \rangle\langle x,y,z \rangle=x+y-2z=0"

"\vec v\cdot\vec n=\langle -1,2,0 \rangle\langle x,y,z\rangle=-x+2y=0"

"\vec w\cdot\vec n=\langle -1,0,1 \rangle\langle x,y,z\rangle=-x+z=0"


"x=2y"

"z=x=2y"

"2y+y-2y=0"

"x=0, y=0, z=0"

Since "x<0," then such vector "\vec n" does not exist.


e)


"\vec u\times \vec v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 0 & -1 & -1 \\
 1 & 0 & -1 \\
\end{vmatrix}"

"=\vec i\begin{vmatrix}
 -1 & -1 \\
 0 & -1
\end{vmatrix}-\vec j\begin{vmatrix}
 0 & -1 \\
 1 & -1
\end{vmatrix}+\vec k\begin{vmatrix}
 0 & -1 \\
 1 & 0
\end{vmatrix}"

"=\vec i-\vec j+\vec k"

"\sqrt{(1)^2+(-1)^2+(1)^2}=\sqrt{3}"


"\vec e=\langle \dfrac{\sqrt{3}}{3}, -\dfrac{\sqrt{3}}{3}, \dfrac{\sqrt{3}}{3}\rangle"